# Clustering Algorithms and RAG Enhancing Semi-Supervised Text Classification with Large LLMs

### v1

@@ 3.1 Clustering Phase
In the initial phase, the emphasis lies on judiciously selecting a limited number of instances from a texture dataset for manual annotation. The dataset, denoted as $X=\{x_{1},x_{2},\ldots,x_{n}\}$, comprises individual documents where each $x_{i}$ symbolizes a sample from the original documents. The objective of this phase is to identify the most effective sampling strategy, $\pi$, under the practical constraint that only a subset $X_{\pi}^{m}$ of $m$ samples can be annotated due to the inherent limitations of human annotation resources. This subset $X_{\pi}^{m}$ of $m$, drawn from the dataset $X$ using the selection method $\pi$, is subsequently assessed by an evaluation metric $Q$, which we elect to be the several metrics that measure the performance of clustering using $X_{\pi}^{m}$. Desbribed in a mathematical way[47], the quest revolves around finding the best selection approach, $\pi^{*}$, which optimizes the chosen performance:
$\pi^{*}=\arg\max_{\pi}Q\left(X_{\pi}^{m},\pi\right)$
Several strategies exist for document data selection, primarily aimed at extracting domain-specific data from more general datasets [48][49]. These methods typically focus on selecting documents in a way that maintains a distribution or style similar to that of an existing target dataset, such as Wikipedia. To attain our specific objective, we utilized unsupervised clustering approaches to identify a greater quantity of smaller and more homogeneous clusters of documents, given the lack of a predefined target distribution. This method is more favorable for identifying representative landmark data that reflect the diversity of the entire dataset, rather than forming a few large clusters consistsing broader categories with a large number of documents. By forming these smaller, more concentrated clusters, we can utilize techniques such as exhaustively analyzing intra-cluster similarity or simply leverage LLMs as the example in Table 3.1 shown to infer and guide the selection of a single, high-quality, representative, most exemplary sample within each cluster. We can then denote the representative landmarks as:
$(X^{m},Y^{m})$
Where $Y^{m}$ are assumed to be annotated by human experts after obtained $X^{m}$.
Table 3.1: Prompt to Choose Representative Landmarks for Each Cluster You have been provided with a set of similar documents, each indexed by a number. Your task is to identify the most representative example from this cluster of documents. Please carefully analyze the given documents and select one document that best captures the common essence and characteristics of the samples. The selection should emphasize the representativeness and relevance of the chosen sample to the category, so that it can serve as a reference for labeling the entire cluster.
Documents in the cluster:
1-. {Document 1}……
2-. {Document 2}……
3-. {Document 3}……
4-. {Document 4}……
$\vdots$
Please choose one document that could best serve as a reference for labeling the entire cluster, and return only the index number of your selection, in format such as [0], [1], etc.
Answer:
3.2 Data Augmentating Phase
During the data augmentation phase, we employ a variety of data augmentation strategies to generate additional training instances. These strategies include using WordNet for word replacement, leveraging LLMs to rewrite documents, employing a combination of CoT reasoning with RAG to predict labels for each unlabeled sample, and directly generating data using RAG. Given the original unlabeled dataset $X$ and the labeled subset $(X^{m},Y^{m})$, our objective during the augmentation phase is to expand the labeled dataset through a range of augmentation techniques denoted as $G$:
$D=\{(x,y):(x,y)\sim G(X,X_{\pi}^{m},Y^{m})\}$
where $x$ stands for an augmented sample, $y$ denotes its corresponding label, and $D$ represents the final augmented dataset used for further model fine-tuning. For augmentation using WordNet and LLMs rewrite methods, only $X_{\pi}^{m}$ is utilized for augmentation purposes, while the corresponding labels $Y^{m}$ remain unchanged. In the case of random replacement using WordNet, a process akin to that described in [26] is adopted, where the WordNet [52] lexical database is used to identify a fixed number of the closest synonyms, and these are then randomly replaced with each other. For LLMs rewrite augmentation, we adopt the prompt strategy as outlined in Table 3.2.
Table 3.2: Prompt For LLMs Document Rewrite *Task Description: Rewrite the following text in English, maintaing the original meaning but using different words and sentence structures. The new version should be clear and concise, and it should not alter the core message of the original text. *Original Text: {The document upon which augmentation will be based} Rewritten Text:
Table 3.3: Prompt For RAG CoT Document Labeling *Task Description: You are provided with a set of similar documents. Your task is to predict the label for the target document, using the labeled document examples as references. *List of Available Labels: {List of available labels, seperated by comma} *Reference Labeled Documents: {A set of labeled landmarks from the nearest five clusters} *Target Document for Prediction: {The document upon which augmentation will be based} *Task: Predict the label for the target document. Please provide your reasoning before asssigning the label. *Format: Thought: [Your thoughts] Label: [Your assigned label] Answer:
For RAG-CoT based labeling of unlabeled samples and generating data using RAG, all $X,X_{\pi}^{m},Y^{m}$ are utilized. Specifically, when employing RAG CoT labeling, we leverage the cluster memberships established during the initial phase to improve our data labeling strategy. For each of the smaller clusters identified, we use LLMs to predict labels for each sample both individually and collectively. Collective predictions allow the LLM to consider the context jointly with other similarly labeled samples from nearby clusters, while CoT judgments help to enhance the accuracy of these predictions. With the details provided in Table 3.3.
Table 3.4: Prompt For RAG Dataset Augmentation *Task Description: You are provided with a set of similar documents, some of which are labeled and others are not. Your task is to generate a sample document based on the primary document, using both the labeled and unlabeled documents as references. *List of Available Labels: {List of available labels, seperated by comma} *Reference Labeled Documents: {A set of labeled landmarks from the nearest five clusters} *Reference Unlabeled Documents: {A set of unlabeled documents within the same cluster} *Primary Document for Augmentation: {The document upon which augmentation will be based} *Task: Using the labeled and unlabeled documents as a guide, create a new document based on the primary document and assign it the appropriate labels from the available list. *Document Format: Content: Label: [Your assigned label] *Generated Example:
@@ 3.2 Data Augmentating Phase
For RAG document generation, we implement a refined approach that leverages both labeled landmarks from similar neighborhood clusters and unlabeled samples within the same clusters, as demonstrated in Table 3.4. For each original document, we identify the top five most probable clusters to which it might belong and select labeled landmarks from these clusters. We then randomly choose three unlabeled samples from the same cluster, combine them with the five labeled landmarks from neighboring clusters, and include the original document itself to construct a diverse prompt. This process does not require incorporating all documents within the cluster, and the random selection of reference documents enhances the diversity of the generated content. As a result, the average and maximum lengths of the prompts were reduced.
3.3 Phase 3: Fine-Tuning using LLMs
In the concluding phase of our project, we leverage the fine-tuning capabilities of pre-trained LLMs to enhance their performance. Fine-tuning LLMs allows us to produce more consistent answers compared to CoT methods and enables smaller models by parameter count to achieve performance comparable to larger models. The prompts designed for training the LLMs are outlined in Table 3.5:
Table 3.5: Prompt For LLM Training Assign tags for the following {subject} Document: {Document} Answer: [label1, label2, … ]
For label prediction tasks, the prompt structure is presented in Table 3.6:
Table 3.6: Prompt For Label Prediction Assign tags for the following {subject} Document: {Document} Answer:
@@ 3.3 Phase 3: Fine-Tuning using LLMs
During the training phase, the model is instructed to output labels in the ‘[label1, label2, …]‘ format. An explicit stop condition is set to ensure that the model begins generating with the opening bracket ‘[‘ and ceases text generation immediately after the closing bracket ‘]‘. Furthermore, to guarantee that the model only produces desired responses, we tokenize the available labels and constrain the model to output only the token IDs associated with these labels. For single-labeled data, the format ‘[label]‘ can be utilized; for multi-labeled data, if the training data are organized by importance from highest to lowest, LLMs will learn to output labels in the sequence ‘[label1, label2, …]‘ reflecting this hierarchy. In hierarchical labeling scenarios, the training data can be structured as ‘[tier1, tier2, …]‘, where ‘tier1‘ labels come from a predefined list of tier1 labels, and ‘tier2‘ labels follow accordingly, thus making the model’s output highly adaptable.
These fine-tuning procedures are conducted within the Llamafactory framework [53]. Our primary goal is to identify the optimal combination of training datasets that yield the best performance metrics when evaluated against test data. We consider various training dataset configurations, such as using landmark data alone, augmenting landmark data with WordNet, LLMs rewriting techniques, or employing RAG. Additionally, we seek to determine the smallest model size that meets performance requirements while minimizing computational resources. The study includes a comparative analysis between fine-tuning results and those obtained by combining RAG and CoT approaches. Specifically, for the CoT and RAG methods, the Qwen2.5-72b model is utilized, whereas for the fine-tuning experiments, the Qwen2.5-0.5B model are employed. This experimental setup enables us to show how different types of training datasets impact performance on specific tasks, offering valuable insights for real-world applications. Notably, all training was conducted within the constraint of an 80 GB GPU memory limit across two A100 GPUs.
4 Metric
4.1 Clustering
4.1.1 Homogeneity
$H(L_{true},L_{pred})=1-\frac{H(L_{pred}\mid L_{true})}{H(L_{true})}$
4.1.2 NMI
$\text{NMI}(U;V)=\frac{I(U;V)}{\text{avg}(\text{H}(U),\text{H}(V))}$
4.1.3 Silhouette Score
$s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$
4.2 Augmentation
4.2.1 Jaccard Similarity
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
5 datasets
@@ 6 Experiment
The initial phase of our experimental methodology involved applying clustering algorithms to the distinct Reuters and Web of Science (WOS) datasets. To validate these methods, we utilized the validation set, leveraging known labels to evaluate clustering performance, whereas we assumed no access to the labels of the training set. From this training set, a carefully selected proportion—300 for the Reuters dataset and 600 for the WoS dataset—was chosen to reveal the true labels. This selection process occurred during Phase 1, where specific data points were identified as landmarks—critical examples particularly valuable for labeling. Specifically, we compared embeddings and TF-IDF models across different cluster sizes and assessed the impact of various clustering methods using metrics such as Normalized Mutual Information (NMI), Homogeneity, and Silhouette Score. Subsequently, we extended our analysis to the training data, reinforcing the findings from the validation phase.
There are a variety of clustering techniques and feature representation methods in our disposal. For the data selection phase, we selected four clustering algorithms and two feature representation strategies that are particularly well-suited for scenarios with large numbers of samples and clusters. The chosen clustering methods include Bisecting K-Means, Balanced Iterative Reducing and Clustering Using Hierarchies (BIRCH)[50], Hierarchical Clustering and GaussianMixture Models. For feature representation, BGE-m3 Embedding[51] and TF-IDF techniques were chosen. The BGE-m3 method embeds up to 8192 tokens into 1024-dimensional vectors. For TF-IDF, we selected severak different max features terms based on frequency scores to address sparsity in the TF-IDF matrix. To evaluate the suitability of these combinations for our specific needs, we measure their performance using several evaluation metrics. Given that we possess the ground truth labels and our goal is to use clustering to identify more representative samples rather than just assigning predicted cluster labels, we are particularly focused on the Homogeneity score, which measures the extent to which each cluster contains only members of a single class. In addition to homogeneity, we also employ the Normalized Mutual Information (NMI) and Silhouette Score to assess the quality of the clustering.
Tf-idf Max Features Homogeneity NMI Silhouette Score
256, TF-IDF 0.7878 0.4784 0.0726
512, TF-idf 0.7929 0.4832 0.0723
1024, TF-idf 0.8104 0.4936 0.0657
2048, TF-idf 0.8003 0.4916 0.0641
4096, TF-idf 0.7895 0.4875 0.0596
-, Embedding 0.8338 0.5123 0.0579
Tf-idf Max Features Homogeneity NMI Silhouette Score
512, TF-idf 0.5058 0.4432 0.0226
1024, TF-idf 0.5455 0.4779 0.0306
2048, TF-idf 0.5573 0.4907 0.0277
4096, TF-idf 0.5547 0.4925 0.0212
8192, TF-idf 0.5453 0.4879 0.0136
-, Embedding 0.6530 0.5677 0.0116
Clustering Method Number of Clusters Homogeneity NMI Silhouette Score Random 100 0.1195 0.0720 / Random 300 0.2679 0.1389 / Random 600 0.4042 0.1934 / Brich 100 0.7299 0.5194 0.0354 Brich 300 0.8147 0.5119 0.0446 Brich 600 0.8593 0.5021 0.0554 Bisecting K-Means 100 0.7101 0.4941 0.0230 Bisecting K-Means 300 0.7916 0.4777 0.0127 Bisecting K-Means 600 0.8473 0.4702 0.0034 Hierarchical 100 0.7467 0.5335 0.0421 Hierarchical 300 0.8338 0.5123 0.0579 Hierarchical 600 0.8839 0.4963 0.0835 Gaussian Mixture 100 0.7513 0.5275 0.0389 Gaussian Mixture 300 0.8127 0.4934 0.0408 Gaussian Mixture 600 0.8765 0.4944 0.0583
From Tables 4 and 5, it can be observed that Hierarchical Clustering and the Gaussian Mixture Model (GMM) outperform other clustering methods in most scenarios. While BIRCH and Bisecting K-Means are algorithms designed for computational efficiency, Hierarchical Clustering and GMM require more processing time, with GMM being the most computationally demanding. In practice, the output of BIRCH is often used as an initial point for hierarchical clustering or GMM. In addition to the encoding method employed, the size of the cluster numbers also impacts the homogeneity and normalized mutual information (NMI) scores. As the number of clusters increases, both homogeneity and NMI scores tend to rise. However, given that the objective of our clustering process is to identify landmark points for labeling, the number of clusters we can manage is constrained by the availability of human resources for labeling. For the Reuters dataset, we selected 300 clusters, whereas for the WOS dataset, we chose 600 clusters.
Clustering Method Number of Clusters Homogeneity NMI Silhouette Score Random 150 0.1565 0.1539 / Random 300 0.2551 0.2345 / Random 600 0.3713 0.3206 / Random 1200 0.4962 0.4044 / Birch 150 0.5141 0.5099 0.0067 Birch 300 0.5785 0.5369 0.0087 Birch 600 0.6420 0.5609 0.0097 Birch 1200 0.7087 0.5859 0.0138 Bisecting K-Means 150 0.4738 0.4685 0.0051 Bisecting K-Means 300 0.5338 0.4943 0.0002 Bisecting K-Means 600 0.5940 0.5168 -0.0077 Bisecting K-Means 1200 0.6642 0.5436 -0.0176 Hierarchical 150 0.5240 0.5202 0.0073 Hierarchical 300 0.5865 0.5440 0.0081 Hierarchical 600 0.6530 0.5677 0.0116 Hierarchical 1200 0.7225 0.5917 0.0190 Gaussian Mixture 150 0.5516 0.5452 0.0206 Gaussian Mixture 300 0.6017 0.5564 0.0177 Gaussian Mixture 600 0.6532 0.5681 0.0130 Gaussian Mixture 1200 0.7033 0.5774 0.0092
We also observed that the Gaussian Mixture Model (GMM) exhibits better performance compared to the Hierarchical Model when the number of clusters is smaller. However, as the size of the clusters increases, the Hierarchical Model surpasses GMM in turns of NMI, Homogeneity and Silhouette Score. An additional advantage of the Gaussian Mixture Model is its probabilistic nature, which enhances the interpretability of cluster membership. GMM provides probabilities for each data point belonging to different clusters, thereby allowing us to identify multiple potential cluster affiliations for each data point. This feature makes GMM particularly suitable for our subsequent step involving rag retrieval. It’s important to note that both Gaussian Mixture Models (GMM) and hierarchical clustering can generate a list of the most likely clusters for each sample. GMM provides a probabilistic framework, assigning a probability density to each sample across all clusters, whereas hierarchical clustering offers a nested structure of subclusters within larger clusters, which can be used to identify the nearest clusters for a given sample.
@@
We then proceed to analyze the training dataset, as illustrated in Figure 1. The resulting clusters exhibit a distribution that is somewhat less skewed than anticipated. Typically, clusters are unevenly distributed, with a majority of elements concentrated within a few large clusters, while many smaller clusters contain only one or two members. For the Reuters dataset, which comprises a total of 300 clusters, the largest cluster contains 100 elements, while the smallest includes just 3 members. The median size of the clusters is 14, and the 95th percentile is at 5 members. The silhouette score for this dataset is 0.0472. For the WOS dataset, consisting of 600 clusters, the largest cluster has 105 elements, and the smallest has only 2 members. The median size of the clusters is 37, and the 95th percentile is at 15 members. The silhouette score for the WOS dataset is 0.1523. Note that metrics requiring true labels are not available, as we assume we only have access to a few labeled training data points in a semi-supervised setting.
[…]
@@
Following the selection and revelation of the true labels of the most representative landmarks from each cluster, we proceeded to utilize all available unlabeled data, along with the labeled landmarks ($X$, $X_{\pi}^{m}$, $Y^{m}$), to generate additional training data. For the RAG generation using prompt as shown in Table 3.4, the average prompt length was 1,836.25 words for the Reuters dataset and 2,260.34 words for the Web of Science (WoS) dataset, with the longest prompts comprising 5,607 and 3,573 words, respectively. For each document, we utilized the prompt to generate three unique samples by varying the randomly selected reference documents. Subsequently, we applied regular expressions to parse out the relevant context and labels from the generated texts. Out of the 16,182 documents generated for the Reuters dataset, 15,473 successfully underwent the regular expression-based extraction of context and labels. Similarly, from the 70,476 documents generated for the WoS dataset, 62,886 documents passed through the extraction process. Additionally, for the rewriting task using large LLMs, as detailed in Table 3.2, we generated ten samples for each of the landmark documents within the datasets. This was conducted with the hyperparameter temperature set to 0.3, yielding a total of 3,000 rewritten documents for the Reuters dataset and 6,000 for the WoS dataset. Since these generated samples did not require label extraction, they all successfully passed the regular expression filtering phase. It is noteworthy that the prompt lengths for the rewriting tasks were significantly shorter compared to those used for RAG generation, averaging 254.95 words for the Reuters dataset and 251.4 words for the WoS dataset.
For the Reuters dataset, we configured the LLMs to produce multiple labels for each document, with individual labels delineated by commas. In contrast, for the WoS dataset, the LLM was instructed to generate a hierarchical labeling structure, beginning with the domain and then specifying the sub-domain, with each level separated by commas. Subsequently, the generated documents were evaluated against their original counterparts using a suite of performance metrics. While the labels for LLMs rewrite are kept the same as the original labels, we investigated the labeling of RAG-generated examples, as depicted in Figure 2. The Reuters dataset originally comprised 90 distinct labels, which were expanded to 443 unique labels through the use of LLMs. This expansion partly results from LLMs generating hallucinatory labels due to lack of fine-tuning for precise label generation. This process utilized 300 labeled landmarks, with 59 labels being covered. Despite this, 93.49% of documents in the augmented dataset still carry labels from the original 90, covering 85 of them. The training set includes 87 out of the original 90 labels. Compared to the original dataset, the augmented version shows increased category sizes. While only 21 categories had over 50 samples originally, now 52 categories exceed this threshold. It is expected that this number will increase further as more samples are augmented. However, there are 9 instances where the number of augmented documents is fewer than the original, often due to rare terms such as "jet," "cotton-oil," "dfl," "cpu," and "naphtha," which appeared only one or two times in the training data. Specifically, the "lei" category, representing the Leading Economic Index, saw no augmentation despite having 8 instances in the training set, likely because LLMs failed to recognize "lei" without context. On average, the augmentation process resulted in 268.14% more documents per category.
For the Web of Science (WoS) dataset which comprising 7 domains and 134 sub-areas, the augmented dataset introduces 11 domains, with 99.90% of the new domains aligned with the original 7. Regarding sub-areas, the augmented dataset generates 2,765 new sub-areas, of which 82.63% remain within the 134 sub-areas of the original WoS dataset, and 133 of the 134 original sub-areas are included. There are 7 cases where the augmented data has fewer samples compared to the original data. Upon examination, the least number of augmented documents, "medical, polycythemia vera" consists of 46, compared to 76 in the training set. The proportionate least is "psychology, attention," with 216 in the training set and only 52 in the augmented data. On average, the augmented data increased the content per category by 178.99%.
@@
Among Gen. Orig. vs Gen. Emb. Sim Among Gen. Orig. vs Gen. Emb. Sim
WordNet 94.57% 97.75% 97.75% 97.13% 98.52% 98.52%
LLM Rewrite 78.36% 47.79% 93.51% 76.32% 46.69% 96.35%
RAG Augment 49.12% 42.31% 92.17% 40.91% 36.08% 90.84%
Table 6: Comparison of different methods on Reuters and WOS datasets.
@@
To evaluate the impact of various text augmentation techniques on the Reuters and WOS datasets, we conducted a comprehensive analysis focusing on the similarity between the original and augmented documents, as well as the diversity among the augmented samples derived from the same source document. Two primary metrics were used to quantify these aspects: Jaccard Similarity, which measures the degree of vocabulary overlap between sets of words, and Embedding Similarity, which assesses semantic similarity. A score closer to 1 indicates a higher degree of similarity. Our findings, summarized in Table 6 and illustrated in Figure 4, show that different augmentation methods yield varying levels of vocabulary overlap and semantic similarity:
@@
  • WordNet Random Synonym Replacement: This method produced the highest vocabulary overlap, with Jaccard Similarity scores of 94.57% for the Reuters dataset and 97.13% for the WOS dataset, based on 10 samples generated from each landmark document, indicating that this technique preserves much of the original text’s structure and vocabulary. Longer documents generally exhibited higher similarity when augmented using WordNet
@@
  • LLMs Rewrite: This method provided a balanced approach, achieving Jaccard Similarity scores of 78.36% for the Reuters dataset and 76.32% for the WOS dataset, based on 10 samples generated from each landmark document. It effectively introduced new variations while maintaining the core essence of the original content.
@@
  • RAG Data Augmentation: This method resulted in the most diverse set of samples, with the lowest Jaccard Similarity scores of 49.12% for the Reuters dataset and 40.91% for the WOS dataset, calculated among the 3 samples generated using the landmark documents and other RAG components. The ability of this method to integrate diverse and randomly selected supporting materials led to significant modifications, thereby enhancing the diversity of the generated content.

### v3

@@ 3 Methods and Datasets
As Figure1 shown, our methodology unfolds in three distinct phases: In the initial phase, the emphasis lies on judiciously selecting a limited number of instances from a texture dataset for manual annotation. The dataset, denoted as $X=\{doc_{1},doc_{2},\ldots,doc_{n}\}$, comprises individual documents where each $doc_{i}$ symbolizes a sample from the original documents. The objective of this phase is to identify the most effective sampling strategy, $\pi$, under the practical constraint that only a subset $X_{\pi}^{m}$ of $m$ samples can be annotated due to the inherent limitations of human annotation resources. This subset $X_{\pi}^{m}$ of $m$, drawn from the dataset $X$ using the selection method $\pi$, is subsequently assessed by an evaluation metric $Q$, which we elect to be the several metrics that measure the performance of clustering using $X_{\pi}^{m}$. Desbribed in a mathematical way[41], the quest revolves around finding the best selection approach, $\pi^{*}=\arg\max_{\pi}Q\left(X_{\pi}^{m},\pi\right)$, which optimizes the chosen performance of metric $Q$. To attain our specific objective, we utilized unsupervised clustering approaches to identify a greater quantity of smaller and more homogeneous clusters of documents, given the lack of a predefined target distribution. This method is more favorable for identifying representative landmark data that reflect the diversity of the entire dataset, rather than forming a few large clusters consistsing broader categories with a large number of documents. By forming these smaller, more concentrated clusters, we then utilize the center of each cluster as our chosen landmarks $X^{m}=\pi^{*}(X)$.
Building on these carefully selected landmarks, we proceed to the annotation phase. During this process, experts not only follow predefined guidelines but also develop their own nuanced standards through iterative refinement. This dynamic approach enhances annotation quality and consistency, ensuring that the synthetic data more accurately reflects real-world scenarios. We denote the representative landmarks as $(X^{m},Y^{m})$, where $Y^{m}$ are annotated by human experts after obtaining $X^{m}$. With these high-quality labeled landmarks in place, we transition into the data augmentation phase. Here, our objective is to expand the labeled dataset through a variety of augmentation strategies. These strategies include leveraging WordNet packages for word substitution, employing LLMs to rewrite documents, and generating data via RAG. In cases of random replacement using WordNet, a method similar to the one described in [20] is adopted. Specifically, the WordNet [42] lexical database is utilized to identify a set number of the closest synonyms, which are then randomly replaced with each other. For LLM-based rewriting augmentation, we follow the prompt strategy detailed in Table C.3. For data augmentation involving WordNet substitutions and LLMs rewriting, only the $X_{\pi}^{m}$ elements undergo augmentation, whereas the corresponding labels $Y^{m}$ remain unaltered.
Task Description Available Tags Retrival "m” Representative Labeled Data with Highest Semantic Similarity Randomly Select "n" Unlabeled Data Points with High Semantic Similarity Prompt Data Output Format Description
Figure 2: RAG Augmentation template for Fine-Tuning Classification LLMs
@@
In our approach to RAG-based document generation, we integrate both labeled landmarks from similar neighborhood clusters and unlabeled samples within those same clusters, as illustrated in Figure 2 and summarized in Table C.4. For each original document, we identify the five most likely clusters it may belong to and select labeled representatives from these clusters. Next, we randomly pick three unlabeled samples from the same cluster, merge them with the five chosen labeled landmarks from neighboring clusters, and combine them with the original document to create a diverse prompt set. This method does not require including all documents within the same cluster; instead, randomly selecting reference documents enhances the diversity of the generated content.
@@
In the concluding phase of our project, we leverage the fine-tuning capabilities of pre-trained LLMs to enhance their performance. Fine-tuning LLMs allows us to produce more consistent answers compared to CoT methods and enables smaller models by parameter count to achieve performance comparable to larger models. The prompts designed for training the LLMs are outlined in Table C.1, while for label prediction tasks, the prompt structure is presented in Table C.2. During the training phase, the model is instructed to output labels in the ‘[label1, label2, …]’ format. An explicit stop condition is implemented so that the model starts generating text with an opening bracket ‘[’ and stops immediately after outputting a closing bracket ‘]’. Furthermore, to guarantee that the model only produces desired responses, we tokenize the available labels and constrain the model to output only the token IDs associated with these labels. For single-labeled data, the format ‘[label]‘ can be utilized; for multi-labeled data, if the training data are organized by importance from highest to lowest, LLMs will learn to output labels in the sequence ‘[label1, label2, …]‘ reflecting this hierarchy. In hierarchical labeling scenarios, the training data can be structured as ‘[tier1, tier2, …]‘, where ‘tier1‘ labels come from a predefined list of tier1 labels, and ‘tier2‘ labels follow accordingly, thus making the model’s output highly adaptable.
The fine-tuning procedures were conducted within the Llamafactory framework [43]. For all experiments, a total of 80GB GPU memory was provided by two A100 GPUs. For tasks that involved augmentation using the LLMs rewrite and RAG methodologies—where no additional training is required—the Qwen2.5-72B model was utilized. Conversely, for the fine-tuning experiments that necessitated iterative training and parameter adjustments, the comparatively smaller Qwen2.5-0.5B model was selected to ensure efficient resource utilization.
@@
4 Metric
4.1 Homogeneity
$H(L_{true},L_{pred})=1-\frac{H(L_{pred}\mid L_{true})}{H(L_{true})}$
4.2 Silhouette Score
$s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$
4.3 Jaccard Similarity
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
@@ 5 Experiment
In the initial phase of our experiment, we applied a variety of clustering algorithms—including Bisecting K-Means, Balanced Iterative Reducing and Clustering using Hierarchies (BIRCH), Hierarchical Clustering, and Gaussian Mixture Models—to both the Reuters and WOS datasets. This was also done to compare the efficacy of embedding techniques against traditional TF-IDF approaches. To validate these methodologies, we used the validation set that included known labels, allowing us to calculate metrics such as Homogeneity, detailed in Appendix A. Following these evaluations, we observed that embedding methods perform better than TF-IDF, and that the Gaussian Mixture Model performs better when the number of clusters is relatively small. Conversely, for a larger number of clusters—such as 1200 or more—Hierarchical Clustering yields superior results. We selected the Gaussian Mixture Model for further analysis because we chose a smaller number of clusters. For feature representation, we employed the BGE-m3[47] Embedding method, which can embed up to 8,192 tokens into a 1,024-dimensional vector.
Figure 3: Homogeneity against cluster size for the Reuters-21578 and WOS-46985 validation datasets, where clusters are assigned by Gaussian Mixture Models (GMM), with homogeneity calculated using true labels.
As illustrated in Figure 3, Homogeneity increases with the number of clusters. It is evident that classifying the WOS dataset poses a greater challenge compared to the Reuters dataset, as indicated by the comparatively lower level of homogeneity achieved by WOS. Additionally, it is observed that randomly assigned cluster membership gradually closes the gap with model-assigned memberships as the number of clusters increases. In scenarios where we encounter a completely new dataset without access to true labels, the silhouette score becomes an valuable metric for assessing clustering quality, as Figure 4 shown. In the Silhouette Score graphs, we see the score begins to decline when the number of clusters exceeds a threshold. For Reuters, this occurs at about 1,800 clusters (33.3% of training data), and for WOS, at around 10,200 clusters (43.4% of training data). However, given that the objective of our clustering process is to identify landmark points for labeling, the number of clusters we can manage is constrained by the availability of human resources for labeling. For the Reuters dataset, we selected 300 clusters, whereas for the WOS dataset, we chose 600 clusters.
Figure 4: Silhouette score against cluster size for the Reuters-21578 and WOS-46985 training datasets, where clusters are assigned by Hierarchical Clustering. Calculating the Silhouette score does not require true labels.
@@
We then proceed to analyze the training dataset, as illustrated in Figure 5. The resulting clusters exhibit a distribution that is somewhat less skewed than anticipated. Typically, clusters are unevenly distributed, with a majority of elements concentrated within a few large clusters, while many smaller clusters contain only one or two members. For the Reuters dataset, which comprises a total of 300 clusters, the largest cluster contains 100 elements, while the smallest includes just 3 members. The median size of the clusters is 14, and the 95th percentile is at 5 members. The silhouette score for this dataset is 0.0472. For the WOS dataset, consisting of 600 clusters, the largest cluster has 105 elements, and the smallest has only 2 members. The median size of the clusters is 37, and the 95th percentile is at 15 members. The silhouette score for the WOS dataset is 0.0152.
Importantly, we assume we did not have access to the labels of the training set. Instead, we strategically selected a proportion of the training data to reveal true labels, identifying key data points as landmarks. To refine the selection of representative samples from each cluster, we implemented a multi-step process. Initially, all samples within each cluster were indexed. Subsequently, LLMs were employed to identify and select the sample that most accurately embodies the core characteristics of its respective cluster. This was achieved by instructing the LLMs to return the index corresponding to the most representative sample, as illustrated by prompts in Appendix C.5. Notably, selecting the sample closest to the center as a representative landmark, calculated through embedding methods, is also an effective approach and the choice between methods can be guided by specific requirements.
@@
Following the selection and revelation of the true labels for the most representative landmarks from each cluster, we proceeded to utilize all available unlabeled data, along with the labeled representatives, to generate additional training data. For RAG generation using the prompt attached in Appendix C.4, we generated three unique samples for each document by varying the randomly selected reference documents. Subsequently, we applied regular expressions to parse out the relevant context and labels from the generated texts. Out of the 16,182 documents generated for the Reuters dataset, 15,473 successfully underwent the regular expression-based extraction of context and labels. Similarly, for the WoS dataset, 62,886 documents passed through the extraction process. Additionally, for the rewriting task using LLMs as prompt detailed in Appendix C.3, we generated ten rewritten samples for each of the landmark documents within the datasets. This was conducted with the hyperparameter temperature set to 0.3, yielding a total of 3,000 rewritten documents for the Reuters dataset and 6,000 for the WoS dataset. Since these generated samples did not require label extraction, they all successfully passed the regular expression filtering phase. Finally, we applied a similar process using WordNet for random replacement, generating an additional 3,000 rewritten documents for the Reuters dataset and 6,000 for the WoS dataset.
@@
Jac. Sim Among Gen. Emb. Sim Orig-Gen Jac. Sim Among Gen. Emb. Sim Orig-Gen
WordNet 94.57% 97.75% 97.13% 98.52%
LLM Rewrite 78.36% 93.51% 76.32% 96.35%
RAG Augment 49.12% 92.17% 40.91% 90.84%
Table 2: Comparison of different augmentation methods on Reuters and WOS datasets
Figure 6: Impact of different Text Augmentation method on Jaccard Similarity (Vocabulary Overlap) and Embedding Similarity (Sementic Similarity) among Generated Samples: The figure illustrates that from traditional WordNet random synonym replacement to LLMs rewrite to RAG augmentation, the vocabulary overlap decrease, indicating the generation of more diverse samples.
@@
To evaluate the impact of various text augmentation techniques on the Reuters and WOS datasets, we conducted a comprehensive analysis focusing on the similarity between the original and augmented documents, as well as the diversity among the augmented samples derived from the same source document. Two primary metrics were used to quantify these aspects: Jaccard Similarity, which measures the degree of vocabulary overlap between sets of words, and Embedding Similarity, which assesses semantic similarity. Our findings, summarized in Table 2 and illustrated in Figure 6, show that different augmentation methods yield varying levels of vocabulary overlap and semantic similarity:
@@
  • WordNet Random Synonym Replacement: This method produced the highest vocabulary overlap, with Jaccard Similarity scores of 94.57% for the Reuters dataset and 97.13% for the WOS dataset, calculated among the 10 samples generated from each landmark document, indicating that this technique preserves much of the original text’s structure and vocabulary.
@@
  • LLMs Rewrite: This method provided a balanced approach, achieving Jaccard Similarity scores of 78.36% for the Reuters dataset and 76.32% for the WOS dataset, calculated among the 10 samples generated from each landmark document. It effectively introduced new variations while maintaining the core essence of the original content.
@@
  • RAG Data Augmentation: This method resulted in the most diverse set of samples, with the lowest Jaccard Similarity scores of 49.12% for the Reuters dataset and 40.91% for the WOS dataset, calculated among the 3 samples generated using the RAG representative documents and other similar random components. The inclusion of varied, randomly chosen supporting documents notably increased content diversity.